OpenStudy (anonymous):

find an equation for the line through (-7,-6) and perpendicular to y=-3/2x+1

hero (hero):

Hint: The the other line you're looking for is perpendicular, then the slope will be \(\dfrac{2}{3}\)

OpenStudy (anonymous):

The slopes should be negative reciprocals!

hero (hero):

If two lines are perpendicular to each other then the product of their slopes will be -1.

hero (hero):

\[-\frac{3}{2} \times \frac{2}{3} =- 1\]

OpenStudy (anonymous):

thus the slopes are negative reciprocals hahaha

OpenStudy (anonymous):

y=mx+b the slope of the 2nd equation is - 3/2

hero (hero):

I think the slope of the second equation is 2/3

OpenStudy (anonymous):

Yup ! then follow this equation ---> y-y1=m(x-x1) to get the next equation x1 and y1 are the given points and m=2/3 since they are negative reciprocals

OpenStudy (anonymous):

That's prolly it. you'll get it after you finalize that equation